OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3). Please help I dont even know where to start.

OpenStudy (anonymous):

Parallel lines have the same slope as the other line.

OpenStudy (anonymous):

Change your line to slope y-intercept format.

OpenStudy (anonymous):

But the line has to pass through the point 9,-3 at the same time. Thats whats confusing me.

OpenStudy (anonymous):

That's okay.

OpenStudy (anonymous):

What is the slope of your line though? Do you recognize that?

OpenStudy (anonymous):

Change your line to y=mx+b format.

OpenStudy (ryan123345):

yes

OpenStudy (ryan123345):

so then it is 3y = -2x + 5

OpenStudy (anonymous):

It would be 3y=-2x+5

OpenStudy (anonymous):

@ryan123345 : let him do the work please.

OpenStudy (anonymous):

Oh. He already posted.

OpenStudy (anonymous):

Okay now divide by 3.

OpenStudy (ryan123345):

ok

OpenStudy (anonymous):

Because we want the format: y=mx+b

OpenStudy (anonymous):

We have 3y.

OpenStudy (anonymous):

Y= - 2/3x+ 3

OpenStudy (anonymous):

Okay but 5/3 is not 3.

OpenStudy (anonymous):

It's just 5/3 .

OpenStudy (anonymous):

So just replace the 3 with 5/3

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

Now do you see what the slope is?

OpenStudy (anonymous):

Ok now i got it. Thank you!

OpenStudy (anonymous):

Okay then. You are welcome :) .